Us the following means and standard deviations: for ACT, µ=21, σ=3 and for SAT, µ=500, σ=100. (a) You take the ACT test and the SAT test. You get a 24 on the ACT and a 660 on the SAT. The college that you apply to only needs one score. Which do you want to send them (that is, which score is better: 24 or 660)? Why?

(b) What is the probability of having an ACT score of 20 or less?

(c) What SAT score do you need to have to be in the top 15% of the population?

(d) What is the probability of scoring between 500 and 650 on the SAT?

(e) What is your percentile rank if you have an ACT of 25.5?
Answer #1

solution Here for Act =21, 6-3 for x=24, 2= *-4 = 24-2) = 1 for SAT, = 500, 2100 for x=660, 2= 66000 - 660-500 = 1.6 here a-